OpenStudy (anonymous):

I need to perform the indicated operations. Round the result to the specified decimal place, when indicated...-3(0.2)-2(3.3) over 30(0.4)^ I have done it for the past hour and cannot come up with the right answer, either that or the answer in the back of the book is not correct. PLEASE help

OpenStudy (anonymous):

Just to make sure we're on the same page:\[\frac{-3\cdot0.2-2\cdot3.3}{30\cdot0.4^1}\]

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

The 30 on the bottom is 30 X 0.4 squaredr

OpenStudy (anonymous):

What do you mean by squaredr? \[30\cdot0.4^2\]?

OpenStudy (anonymous):

yes sorry keyboard issues

OpenStudy (anonymous):

Ok, well 0.4^2=0.16 and 0.16*30=4.8 So we have 4.8 in the denominator

OpenStudy (anonymous):

-3*0.2 -2*3.3 = -0.6 - 6.6 = -7.2 So -7.2 on top.

OpenStudy (anonymous):

which gives -7.2/4.8=-1.5